about how many times might you expect to get a number greater than 4 when you toss a number cube 400 times?
Answer by ewatrrr(24785)


Hi
about how many times might you expect to get a number greater than 4 when you toss a number cube 400 times
5,6 are the two, out of the six outcomes, > 4
P(number > 4) = 2/6 = 1/3 OR 33.33% of the time it would be expected to happen.
.3333*400 = 133 times might expect th enumber to be >4